What acute angle does a line with slope -3/2 make with a vertical line? Write answer in decimal
form.

Answer by ikleyn(52803)

.
What acute angle does a line with slope -3/2 make with a vertical line? Write answer in decimal form.
~~~~~~~~~~~~~~


arctan(-3/2) = -56.31 degrees  (rounded).


This angle is in QIV.  The corresponding angle in QII is  -56.31 + 180 = 123.69 degrees.


Hence, the acute angle with vertical axis is 123.69 - 90 = 33.69 degrees.    ANSWER

Solved.